OpenStudy (anonymous):

If x = -3, then x^2-7x + 10 equals

OpenStudy (anonymous):

\[(-3)^2-7(-3)+10=9+21+10=40\] you can use this solution too. \[(x-2)(x-5)=(-5)\times(-8)=40\]

OpenStudy (anonymous):

If you look back to our last session, you'll see I wrote out the steps to the solution, but @amirreza1870 is exactly right as well.

OpenStudy (anonymous):

= (-3)^2-7(-3)+10 = 9+21+10 =40

OpenStudy (usukidoll):

errr you just plug -3 into the equation

OpenStudy (anonymous):

thanks so ok x^2 is -3*-3 right then its going to be -7*-3

OpenStudy (anonymous):

Solve 3x - 1 = 10

OpenStudy (anonymous):

is x=3 right

OpenStudy (usukidoll):

-_-

OpenStudy (usukidoll):

+ 1 on both sides.. divide by 3 x = 11/3